Calculate the number of miles of calcium chloride that can be obtained from 25g of limestone in the presence of excess hydrogen chloride
1
Expert's answer
2020-11-20T10:03:07-0500

CaCO3+2HCl=CaCl2+H2CO3

n (CaCO3) = n (CaCl2)

m (CaCO3)=25 g

M (CaCO3)=100.1 g/mol

n (CaCO3)=m/M = 25/100.1 = 0.25 mol

n (CaCl2)=n (CaCO3) = 0.25 mol